There is 20 maggots in 100g of mushrooms. How many maggots are in 350g of mushrooms?
Answer by stanbon(75887)

There is 20 maggots in 100g of mushrooms. How many maggots are in 350g of mushrooms?
-----
Use a proportion:
x/350g = 20/100g
---
x = 350(20/100) = 350(1/5) = 70